I already know the graph of f for -1≤x≤5. Now I need to find the value of 'p' in the equation: f(x)=(x-p)^2-3.
Thank you!

Answer by Theo(13342)

f(x) = (x-p)^2 - 3.

if you know what f(x) is, then you can solve for p.

for example, if f(3) = 100, then f(x) = (x-p)^2 - 3 becomes f(3) = (3-p)^2 - 3 which becomes 100 = (3-p)^2 - 3

add 3 to both sides to get 103 = (3-p)^2

take the squaare root of both sides to get sqrt(103) = 3 - p

subtract 3 from both sides to get sqrt(103) - 3 = - p

multiply both sides by -1 to get 3 - sqrt(103) = p

solve for p to get p = 3 - sqrt(103) = -7.148891565

when x = 3 and p = -7.148891565, the formula becomes 100 = (3 - (-7.148891565))^2 - 3 which becomes 100 = (3 + 7.148891565)^2 - 3 which becomes 100 = 10.148891565^2 - 3 which becomes 100 = 100.

this confirms the solution is correct, given the assumptions.